# Analysis of Fourier Neural Operators via Effective Field Theory

### v2

@@ 5.2 Susceptibility Prediction
This section validates the parallel ($\chi_{\parallel}$) and perpendicular ($\chi_{\perp}$) susceptibilities predicted by our theory. We fix the architecture to truncation $k_{\max}=32$ and channel width $32$, with inputs drawn from the same Gaussian field as above, and read out the first Fourier layer. For the parallel tests we add a perturbation of size $\epsilon=10^{-2}$ along the theory-specified data direction and compute $\chi_{\parallel}(f)$ via Eqs. 25, 29 and 32 averaging over 100 seeds. For the perpendicular tests we inject an orthogonal perturbation of size $\epsilon=10^{-5}$, integrate the resulting power over all frequencies following Eqs. 26, 30 and 33 and again average over 100 runs. We evaluate four activation settings: quadratic, cubic, ReLU, and a ResNet variant with a cubic activation and residual gain $\gamma\in\{0.25,0.75\}$. Except for the $C_{R}(f)\propto 1/|f|$ ablation, the weight-variance profile $C_{R}(f)$ is a step function that vanishes above $k_{\max}$; for the parallel tests we fix $C_{R}=1$, while for the perpendicular tests we sweep $C_{R}\in[0.80,1.20]$. Because non-ResNet models have zero kernel beyond $k_{\max}$, we show their curves only up to $k_{\max}$; the ResNet retains post-cutoff energy, so we plot it up to $2k_{\max}$. Figure 1 shows close agreement between theory and measurement: empirical susceptibilities lie within one standard deviation of the predicted curves across activations, perturbation types, and depths.
@@ 6 Conclusion
We analyze neural operators directly in function space, characterizing their frequency domain kernels and susceptibilities without resorting to discretization. The theory reveals that convolution in Fourier Neural Operators induces frequency coupling: even after spectral truncation, a reduced kernel retains energy at higher modes. The resulting behavior depends systematically on activation class—analytic versus non-analytic (with singular features)—and the predictions are borne out experimentally: the kernel evolution recursion matches measured kernels across depth, and both parallel and perpendicular susceptibilities agree with theory within empirical one-sigma bands.
These findings will provide practical criteria for selecting initialization hyperparameters and for anticipating spectral transfer as architectures deepen or widen. Promising next steps are to turn these criteria into concrete initialization/tuning procedures, broaden the analysis beyond the analytic, scale invariant, and residual cases, track the ensemble distribution through training, and investigate higher-order correlations.
@@ Appendix A FCN Effective-Theory Background
Now, the distribution of $l$-th layer pre-activation conditioned on l1l-1-th layer is also following Gaussian distribution:
@@
It can be easily checked that $2k$-point connected correlators is $\mathcal{O}\Big{(}\frac{1}{{n^{(l)}}^{k}}\Big{)}$. So, for large enough widths, we can consider the neural network ensembles as a Gaussian processes. As widths go to infinity, the $2$-point correlations go to some fixed kernel let denote this kernel as $K^{(l)}_{\alpha_{1}\alpha_{2}}$. According to Roberts et al. (2022), for infinite width neural networks, the running of kernels and four-points vertices are described as follows:
@@ Appendix B Scaling Law and Four-Point Vertex
Four-point vertex We define the fluctuation of metric $\Delta\widetilde{\mathcal{G}^{(l)}}$ as follows equation:
@@
and define four-point vertex $\mathcal{V}^{(l)}\{(\mathbf{u}_{1},\mathbf{u}_{2}),(\mathbf{u}_{3},\mathbf{u}_{4})\}$ which is scaled variance of fluctuation $\Delta\widetilde{\mathcal{G}^{(l)}}$. Using this quantity, we can calculate four-point correlators perturbatively. Eq. 19 shows the expansion of four point vertex which is composed of four-point connected correlators of (l1)(l-1)-th layer activations.
@@ Appendix D Proofs for Theorem 1
$\begin{split}&\tilde{\chi}_{\perp}(f,f^{\prime})=\|\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\})\|_{\mathcal{K}^{(l)}\{u_{0}\}}\\ &=\sum_{i}\frac{1}{n^{(l)}}\|\sigma_{1}+\sigma_{2}R^{(l)}_{ij}(f)\hat{u}_{j}(f)+\frac{\sigma_{3}}{2!}\Big{(}(R^{(l)}_{ij}\hat{u}_{j})*(R^{(l)}_{ij^{\prime}}\hat{u}_{j^{\prime}})\Big{)}(f)+\cdots\|_{\mathcal{K}^{(l)}\{u_{0}\}}\\ &=\delta(f-f^{\prime})\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\sum_{n_{k,1}+\dots+n_{k,l}=k-1}\frac{(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f).\end{split}$
@@ Appendix E Proofs for Theorem 2
To compute the covariance value of absolute of random variables, we need following lemmas from Heydenreich & Hofstad (2009), Vleck (1943).
@@
$\begin{split}\mathbb{E}[I(x)]&=\frac{1}{\sqrt{2\pi}}V^{\frac{1}{2}}\quad\text{for all }x\in\mathbb{R},\\ \mathbb{E}[H(x)]&=\frac{1}{2}\quad\text{for all }x\in\mathbb{R},\\ \mathbb{E}[I(x)\overline{I}(x^{\prime})]&=\frac{V}{4\pi}(2\sqrt{1-\rho^{2}}+\rho(\pi+2\arcsin{\rho}),\\ \mathbb{E}[H(x)\overline{H}(x^{\prime})]&=\frac{1}{4}+\frac{1}{2\pi}\arcsin{\rho},\\ \mathbb{E}[\overline{R}(f^{\prime})\overline{U}(f^{\prime})\mathcal{F}(I(x))]&=\frac{1}{2}\delta(f-f^{\prime})K(f,f^{\prime})C(f)\end{split}$
@@
Now, consider the Fourier transform $\mathcal{F}(\mathfrak{S}^{(l)}\{\mathbf{u}\})$ in the case of scale invariant activation. We have
@@ Appendix F Derivation of Theorem 3
$\begin{split}&\langle\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\}+\tilde{\gamma}),\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\}+\tilde{\gamma})\rangle_{\mathcal{K}^{(l)}\{u_{0},u_{0}\}}\\ &=\langle\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\}),\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\})\rangle_{\mathcal{K}^{(l)}\{u_{0},u_{0}\}}+2\langle\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\}),\tilde{\gamma}\rangle_{\mathcal{K}^{(l)}\{u_{0},u_{0}\}}+\gamma\delta(f)\delta(f^{\prime})\\ &=\Big{(}\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\sum_{n_{k,1}+\dots+n_{k,l}=k-1}\frac{(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)\Big{)}\\ &+2\tilde{\gamma}\sigma_{1}\delta(f)\delta(f^{\prime})+\gamma\delta(f)\delta(f^{\prime}).\end{split}$

### v3

@@ 5.3 Susceptibility Prediction
We validate the parallel ($\chi_{\parallel}$) and perpendicular ($\chi_{\perp}$) susceptibilities predicted by our theory by comparing the theoretical curves with empirical measurements under random initialization. We fix the architecture to truncation $k_{\max}=32$ and width $n=32$, draw inputs from the same Gaussian random field as in Section 5.2, and read out the first Fourier layer. For the parallel experiment, we add a perturbation of amplitude $\epsilon_{\parallel}=10^{-2}$ along the theory-specified data direction and compute $\chi_{\parallel}(f)$ using Eqs. 25, 29 and 32, averaging over $N=100$ seeds. For the perpendicular experiment, we inject an orthogonal perturbation of amplitude $\epsilon_{\perp}=10^{-5}$ and compute $\chi_{\perp}$ by integrating the resulting power over frequencies as in Eqs. 26, 30 and 33, again averaging over $N=100$ runs.
Activation settings and spectral profiles.
We test four settings: quadratic, cubic, ReLU, and a ResNet variant with cubic activation and residual gain $\gamma\in\{0.25,0.75\}$. Except for the decaying-spectrum ablation ($C_{R}(f)\propto 1/|f|$ with truncation), we use a step profile $C_{R}(f)=\mathbf{1}_{\{|f|\leq k_{\max}\}}$. For the parallel tests we fix the in-band scale to $C_{R}=1$, while for the perpendicular tests we sweep $C_{R}\in[0.80,1.20]$ to probe stability around the criticality condition.
Results.
Because non-residual models have negligible energy beyond $k_{\max}$, we display their curves only up to $k_{\max}$, whereas residual models retain post-cutoff energy and are therefore plotted up to $2k_{\max}$. In Fig. 10 and 11, markers denote empirical measurements and solid curves denote theory; error bars indicate $\pm 1$ standard deviation over $N=100$ seeds. Across all activations, perturbation types, and tested depths, the empirical susceptibilities closely track the theoretical predictions: measurements remain within one standard deviation of the predicted curves throughout the retained band, and, where applicable, the residual model continues to match accurately beyond $k_{\max}$.
As shown in Fig. 10, under the step profile $C_{R}(f)=\mathbf{1}_{\{|f|\leq k_{\max}\}}$ we observe a pronounced bump in the spectrum near the truncation boundary, in contrast to the smoother behavior under decaying profiles. We attribute this to an artifact induced by the discontinuity of the step profile: the sharp cutoff introduces oscillatory components and amplifies boundary effects when propagated through the convolutional structure of the kernel recursion. This observation suggests a practical refinement for initialization: choosing a smoothly decaying $C_{R}$ near $k_{\max}$ may yield a more stable neural operator by mitigating truncation-edge artifacts.
In Fig. 11, the empirical susceptibilities are systematically slightly larger than the theoretical predictions. A plausible explanation is that our prediction retains only the leading two-point statistics (kernel-level theory), while finite-width corrections—controlled by higher-order connected correlators such as the four-point vertex—are neglected. Incorporating these higher-order terms should improve quantitative accuracy and may account for the observed upward bias.
5.4 Computational Complexity and Scalability
Cost and memory under the baseline (A3).
Under Assumption (A3) (frequency-diagonal spectral covariance), the layer kernel remains diagonal in frequency, so it is represented by a single spectrum $\mathcal{K}^{(l)}(f)$ rather than a dense $\mathcal{K}^{(l)}(f,f^{\prime})$. On a $d$-dimensional periodic grid with $N_{x}^{d}$ points and a truncation radius $k_{\max}$, the number of retained modes is $M\asymp(2k_{\max}+1)^{d}$; storing $\mathcal{K}^{(l)}$ costs $\mathcal{O}(M)$ memory. This diagonal structure is the main reason the recursions are tractable in practical FNO settings.
Analytic activations (convolution powers).
For analytic $\sigma$, the recursion in Theorem 1 involves convolution powers of $\mathcal{H}^{(l)}$ up to some order $K$. Each $k$-fold convolution can be computed via FFTs, since convolution in the frequency domain corresponds to pointwise multiplication in the spatial domain. Moreover, because there are a total of $2^{(K-1)}$ combinations at each order, summing all of them yields a per-layer cost
$\mathcal{O}\!\big(K2^{K}M\log M\big)$
for fixed $K$. In practice, the effective order can be truncated: the $k$-th term is suppressed by $(n^{(l)})^{-(k-1)}$ (Theorem 1), so for wide layers only low orders contribute above numerical noise, and we truncate the series once the remaining tail falls below a target tolerance.
Scale-invariant activations (ReLU-type).
For Theorem 2, the update can be implemented by computing the spatial correlation $\rho^{(l)}(x-x^{\prime})$ from $\mathcal{H}^{(l)}$ via an inverse FFT, applying the closed-form nonlinearity map in the spatial domain, and transforming back by FFT. This yields a per-layer complexity $\mathcal{O}(M\log M)$ in any dimension $d$ (up to constant factors).
Four-point vertex.
Exact evaluation of the four-point vertex is indeed expensive. However, in the wide-network regime the vertex is $\mathcal{O}(1/n)$ and the kernel recursion closes at leading order, which is the regime emphasized in this work. When finite-width corrections are desired, a practical option is Monte-Carlo estimation across initialization seeds: compute the sample covariance of the metric fluctuation $\Delta\widetilde{\mathcal{G}}^{(l)}$ at selected frequencies, or estimate only the contracted quantities that enter susceptibilities using stochastic trace estimators. Both approaches avoid forming a dense four-frequency tensor and keep the cost proportional to the number of probed modes.
5.5 Training Dynamics
Stability-criterion calibration.
In this section, based on the theoretical results derived above, we propose an algorithm for selecting hyperparameters that yield a stable network configuration (i.e. satisfying our criticality criterion). Unlike the FCN case, the spectral profile is a function, which admits infinitely many possible candidate algorithms. Here, we present one concrete choice among them.
From the empirical results in Section 5.3, we observe that when the profile contains an abrupt discontinuity, the parallel susceptibility can exhibit spiky behavior. To avoid this issue, we assume a profile that decays smoothly up to the truncation frequency $k_{\max}$. As a representative example, we consider the following class of profiles:
$\Big\{\frac{s}{1+\exp(-c(\tilde{k}_{\text{decaying}}-|f|))}1_{\{|f|<k_{\max}\}},s>0,c>0\Big\}.$
This function class is parameterized by a scale factor $s$, a steepness parameter $c$ controlling how rapidly the profile decays, and a fixed decay onset $k_{\text{decaying}}$ ($\tilde{k}_{\text{decaying}}=\frac{k_{\text{decaying}}+k_{\max}}{2}$). In our algorithm, $k_{\text{decaying}}$ effectively replaces the role of the previous $k_{\max}$: the model is treated as handling frequencies up to $k_{\text{decaying}}$, and the parallel susceptibility is enforced to be 1 only up to $k_{\text{decaying}}$. In contrast, the perpendicular susceptibility is considered up to $k_{\max}$.
Algorithm 1 describes the procedure for finding hyperparameters $(s^{\star},c^{\star})$ that satisfy the susceptibility criteria, given the data distribution, $k_{\text{decaying}}$, $k_{\max}$, and architectural parameters such as width. Based on this algorithm, we conduct the experiment presented in the following paragraph.
Algorithm 1 Calibration of logistic spectral profile for parallel/perpendicular criticality
1:Dataset (or input distribution) $\mathcal{D}$ over functions $u$, cutoffs $k_{\text{decaying}},k_{\max}$, width $n$, depth $L$, activation class (analytic / ReLU-type / residual), series truncation $K_{\mathrm{trunc}}$ (if analytic), weights $w(f)\geq 0$ for parallel fitting.
2:Profile $C_{R}(f;s,c)=s\,b_{c}(f)$, $b_{c}(f)=\big(1+\exp(-c\,g(f))\big)^{-1}$, $g(f)=\tilde{k}_{\text{decaying}}-|f|$. $\triangleright$ $\tilde{k}_{\text{decaying}}=\frac{k_{\text{decaying}}+k_{\max}}{2}$
3:Calibrated $(s^{\star},c^{\star})$.
4:Sample $\{u^{(m)}\}_{m=1}^{M}\sim\mathcal{D}$ and compute $\hat{u}^{(m)}(f)$ for $f\in\mathcal{F}$.
5:Compute input-spectrum statistics (mean included):
$\mu(f)\leftarrow\frac{1}{M}\sum_{m=1}^{M}\hat{u}^{(m)}(f),\qquad S(f)\leftarrow\frac{1}{M}\sum_{m=1}^{M}|\hat{u}^{(m)}(f)|^{2}.$
$\triangleright$ $S(f)=\mathrm{Var}(\hat{u}(f))+|\mu(f)|^{2}$
6:Define the target layer index $\ell$ at which the criterion is enforced (e.g., first Fourier layer).
7:Initialize search range $c\in[c_{\min},c_{\max}]$ and $s\in[s_{\min},s_{\max}]$.
8:for candidate $c$ in a 1D search scheme (grid / golden-section) do
9:  Set $b_{c}(f)\leftarrow(1+\exp(-c\,g(f)))^{-1}$. $\triangleright$ (A) Inner solve: enforce the perpendicular criticality as a scalar root-finding in $s$
10:  Define the function
$\Phi_{\perp}(s;c)\;:=\;\sum_{f\in[0,f_{\text{max}}]}C_{R}(f;s,c)\,\tilde{\chi}_{\perp}^{(\ell)}(f;s,c)\,\Delta f\;-\;1,$
where $\tilde{\chi}_{\perp}^{(\ell)}$ is computed from the theory (Cor. 1/2/3) using the kernel recursion up to layer $\ell_{0}$, with $C_{R}(\cdot)=s\,b_{c}(\cdot)$ and (if analytic) truncation $K_{\mathrm{trunc}}$. $\Delta f$ is the spacing of frequency domain.
11:  Solve $\Phi_{\perp}(s;c)=0$ for $s$ via bisection (or safeguarded Newton), yielding $s(c)$. $\triangleright$ (B) Evaluate parallel deviation under the perpendicular-calibrated $s(c)$
12:  Compute $\chi_{\parallel}^{(\ell)}(f;s(c),c)$ for all $f\in[0,f_{\text{decaying}}]$ from the corresponding susceptibility formula (e.g. Eqs. 25, 29 and 32).
13:  Compute a parallel mismatch score, e.g.,
$E_{\parallel}(c)\;:=\;\sum_{f\in[0,f_{\text{decaying}}]}w(f)\big(\chi_{\parallel}^{(\ell)}(f;s(c),c)-1\big)^{2}\quad\text{or}\quad E_{\parallel}^{\infty}(c)\;:=\;\max_{f\in[0,f_{\text{decaying}}]}\big|\chi_{\parallel}^{(\ell)}(f;s(c),c)-1\big|.$
14:end for
15:Choose $c^{\star}\leftarrow\arg\min_{c}E_{\parallel}(c)$ (or $\arg\min_{c}E_{\parallel}^{\infty}(c)$), and set $s^{\star}\leftarrow s(c^{\star})$.
16:return $(s^{\star},c^{\star})$.
Scope of the theory (early-training regime).
Our kernel and susceptibility-based analysis is primarily valid at initialization and during the early stage of training, as discussed in our previous work (Kim and Kang (2024b)). As optimization proceeds, higher-order correlation functions (e.g., the 4-point vertex) can grow, and the effective kernel may deviate from its initial form, leading to more complex spectral dynamics beyond the present theory. Accordingly, Algorithm 1 should be interpreted as an initialization (or early-training) calibration that enforces criticality at the target layer $\ell$, rather than a guarantee that the criterion remains exactly matched throughout training.
Benchmark results under criterion-matched initialization.
We compared the training loss and test error of a vanilla FNO and an FNO equipped with our proposed algorithm on the 1D Burgers equation. We used the PDEBench (Takamoto et al. (2022)) 1D Burgers dataset with viscosity $\nu=0.001$ for training. The input corresponds to $T=0$ and the target to $T=0.1$, and we used the full positional grid of 1024 points provided in the original dataset. The training set consists of 1000 samples.
For the model, we used a 1D FNO with $k_{\max}=128$, width =32=32, and 5 Fourier layers, with $\tanh$ as the activation function. For the model with our algorithm, we considered the analytic-activation setting with $K_{\mathrm{trunc}}=3$ and $k_{\mathrm{decaying}}=64$, and searched over $s\in[0,2]$ and $c\in[0.01,3]$, using the $\ell^{\infty}$ parallel mismatch score.
For training, we used the Adam optimizer with learning rate $10^{-3}$ and weight decay $3\cdot 10^{-3}$. The batch size was set to 50, and we trained for 1000 epochs. The training objective was the relative $L^{2}$ error. The training-loss curves are shown in Fig. 12, and the test performance is reported in Table 1.
As shown in Fig. 12, the FNO initialized via our criterion-matched calibration exhibits markedly more stable training dynamics: the shaded variability band across seeds is tightly concentrated around the mean throughout training, indicating low sensitivity to random initialization. In addition, the calibrated model converges faster and achieves a lower final training loss than the vanilla baseline.
Table 1 reports the mean and standard deviation of the relative $L^{2}$ error evaluated on an independent test set of 1000 samples drawn from PDEBench. The calibrated model attains a lower test loss than the vanilla FNO, demonstrating that the proposed initialization calibration improves not only early-training stability but also downstream generalization performance.
Figure 12: Training loss dynamics (mean $\pm 1$ std). Curves show the mean training loss across 5 repeated random initializations, with shaded regions indicating $\pm 1$ standard deviations. The calibrated model (ours) exhibits more stable optimization and faster reduction in training loss compared to the vanilla FNO baseline.
Model Mean test loss Std. dev.
Vanilla FNO 0.451 $3.32\times 10^{-5}$
Calibrated (Ours) 0.432 $1.24\times 10^{-4}$
Table 1: Test set loss after training (mean and standard deviation over 5 random initializations).
@@ 6 Conclusion
We analyze neural operators directly in function space, characterizing their frequency-domain kernels and susceptibilities without resorting to discretization. The theory reveals that convolution in Fourier Neural Operators induces frequency coupling: even after spectral truncation, a reduced kernel retains energy at higher modes. The resulting behavior depends systematically on activation class—analytic versus non-analytic (with singular features)—and the predictions are borne out experimentally: the kernel evolution recursion matches measured kernels across depth, and both parallel and perpendicular susceptibilities agree with theory within empirical one-sigma bands.
Beyond early-training predictions, we also show that the criticality theory leads to actionable improvements in practical training. Concretely, our criterion-matched initialization (calibration) produces markedly more stable optimization across random seeds, faster convergence, and lower achieved loss; on a standard PDEBench Burgers benchmark it also improves test performance relative to a vanilla FNO baseline, indicating better downstream generalization.
These findings provide practical criteria for selecting initialization hyperparameters and for anticipating spectral transfer as architectures deepen or widen. Promising next steps are to further develop these criteria into automated initialization/tuning procedures, broaden the analysis beyond the analytic, scale-invariant, and residual cases, track the ensemble distribution through training, and investigate higher-order correlations.
@@ Appendix A FCN Effective-Theory Background
Now, the distribution of $l+1$-th layer pre-activation conditioned on $l$-th layer is also following Gaussian distribution:
@@
It can be easily checked that $2k$-point connected correlators is $\mathcal{O}\Big(\frac{1}{{n^{(l)}}^{k-1}}\Big)$. So, for large enough widths, we can consider the neural network ensembles as a Gaussian processes. As widths go to infinity, the $2$-point correlations go to some fixed kernel let denote this kernel as $K^{(l)}_{\alpha_{1}\alpha_{2}}$. According to Roberts et al. (2022), for infinite width neural networks, the running of kernels and four-points vertices are described as follows:
@@ Appendix B Scaling Law and Four-Point Vertex
[…]
@@
and define four-point vertex $\mathcal{V}^{(l)}\{(\mathbf{u}_{1},\mathbf{u}_{2}),(\mathbf{u}_{3},\mathbf{u}_{4})\}$ which is scaled variance of fluctuation $\Delta\widetilde{\mathcal{G}^{(l)}}$. Using this quantity, we can calculate four-point correlators perturbatively. Eq. 19 shows the expansion of four-point vertex which is composed of four-point connected correlators of (l1)(l-1)-th layer activations.
@@ Appendix D Proofs for Theorem 1
$\begin{split}&\tilde{\chi}_{\perp}(f,f^{\prime})=\|\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\})\|_{\mathcal{K}^{(l)}\{u_{0}\}}(f,f^{\prime})\\ &=\sum_{i}\frac{1}{n^{(l)}}\|\sigma_{1}+\sigma_{2}R^{(l)}_{ij}(f)\hat{u}_{j}(f)+\frac{\sigma_{3}}{2!}\Big((R^{(l)}_{ij}\hat{u}_{j})*(R^{(l)}_{ij^{\prime}}\hat{u}_{j^{\prime}})\Big)(f)+\cdots\|_{\mathcal{K}^{(l)}\{u_{0}\}}\\ &=\delta(f-f^{\prime})\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\sum_{n_{k,1}+\dots+n_{k,l}=k-1}\frac{(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f).\end{split}$
@@ Appendix E Proofs for Theorem 2
To compute the covariance value of absolute of random variables, we need following lemmas from Heydenreich and Hofstad (2009), Vleck (1943).
@@
$\begin{split}\mathbb{E}[I(x)]&=\frac{1}{\sqrt{2\pi}}V^{\frac{1}{2}}\quad\text{for all }x\in\mathbb{R},\\ \mathbb{E}[H(x)]&=\frac{1}{2}\quad\text{for all }x\in\mathbb{R},\\ \mathbb{E}[I(x)\overline{I}(x^{\prime})]&=\frac{V}{4\pi}\Big(2\sqrt{1-\rho^{2}}+\rho(\pi+2\arcsin{\rho})\Big),\\ \mathbb{E}[H(x)\overline{H}(x^{\prime})]&=\frac{1}{4}+\frac{1}{2\pi}\arcsin{\rho},\\ \mathbb{E}[\overline{R}(f^{\prime})\overline{U}(f^{\prime})\mathcal{F}(I(x))]&=\frac{1}{2}\delta(f-f^{\prime})K(f,f^{\prime})C(f)\end{split}$
@@
Now, consider the Fourier transform $\mathcal{F}(\mathfrak{S}^{(l)}\{\mathbf{u}\})$ in the case of scale-invariant activation. We have
@@ Appendix F Derivation of Theorem 3
$\begin{split}&\langle\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\}+\tilde{\gamma}\mathbf{1}),\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\}+\tilde{\gamma}\mathbf{1})\rangle_{\mathcal{K}^{(l)}\{u_{0},u_{0}\}}\\ &=\langle\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\}),\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\})\rangle_{\mathcal{K}^{(l)}\{u_{0},u_{0}\}}+2\langle\mathcal{F}(\mathfrak{S}^{\prime(l)}\{\mathbf{u_{0}}\}),\tilde{\gamma}\mathbf{1}\rangle_{\mathcal{K}^{(l)}\{u_{0},u_{0}\}}+\gamma\delta(f)\delta(f^{\prime})\\ &=\delta(f-f^{\prime})\Big(\sum_{k=1}^{\infty}\frac{\sigma_{k}^{2}}{(n^{(l)})^{k-1}}\sum_{n_{k,1}+\dots+n_{k,l}=k-1}\frac{(n_{k,1})!\dots(n_{k,l})!}{(k-1)!}(\mathcal{H}^{(l)})^{*n_{k,1}}\cdots(\mathcal{H}^{(l)})^{*n_{k,l}}(f)\Big)\\ &+2\tilde{\gamma}\sigma_{1}\delta(f)\delta(f^{\prime})+\gamma\delta(f)\delta(f^{\prime}).\end{split}$
Here $\mathbf{1}=(1,\dots,1)^{T}\in\mathbb{R}^{n^{(l)}}$.